Woman in the military ranks. the percentage of woman in the total active military duty force has been steadily increasing. the number N of woman in the active duty force t years after 1971 is approximated by N=12,197.8t+44,000. How many women were in the military in 1971(t=0)? In 1981(t=10)? In 1990(t=19)?
Answer by stanbon(75887)

Woman in the military ranks. the percentage of woman in the total active military duty force has been steadily increasing. the number N of woman in the active duty force t years after 1971 is approximated by N=12,197.8t+44,000.
How many women were in the military in 1971(t=0)?
Ans: 44,000
------------------
In 1981(t=10)?
N(10) = 12197.8*10 + 44000
N(10) = 121978 + 44000
N(10) = 165978
-------------------
In 1990(t=19)
N(19) = 12197.8*19 + 44000
N(19) = 275758 when rounded down
------------
Cheers,
Stan H.
-------